OpenStudy (anonymous):

The area of a window is 18ft2.<-(square feet) the length of the window is two times the width of the window. What are the dimensions of the window?

OpenStudy (anonymous):

\[18ft^{2}\]

OpenStudy (anonymous):

We know that l*w is the area, so if we substitute l for two times w, we get 2w+w=18, and you combine the like terms to get 3w=18. balance both sides out to get 3w/w=18/3, so basically divide 18 by 3 to get the width. Hope this helps :)

OpenStudy (anonymous):

wait a sec I made a mistake :( it's actually 3w/3=18/3, not 3w/w=18/3.

OpenStudy (anonymous):

Okay..

OpenStudy (anonymous):

you know that two times width is length, right?

OpenStudy (anonymous):

so instead of l*w=18, replace l with 2w

OpenStudy (anonymous):

So each side is 3 ft is the answer?... @jchiang